In order to determine that the precision of leveling equipment is appropriate for the intended measuring task, two calibration tests were carried out with two samples of measurements by the same instrument but different observers. The results of the tests are as follows:
Test 1:
Computed standard deviation of instrument = ± 2.5 mm;
number of degrees of freedom = 24.
Test 2:
Computed standard deviation of instrument = ± 2.0 mm;
number of degrees of freedom = 15.
Construct the 90% confidence interval for the ratio of the two population variances associated with the two tests.
Answer #1

Confidence interval for the ratio of the two population variances is given by,

Lower bound =    Fo/2 ;  F(α/2) follows f distribution with d.f1 = n1-1 and d.f2 =n2-1

Upper bound =  2 F1-(a/2) ; F1-(α/2) follows f distribution with d.f1 = n1-1 and d.f2 =n2-1

We are given S1 = 2.5 and S2 = 2 ; therefore  S12 = 6.25 and S22 = 4

Therefore  S12 /  S22 = 6.25/4 = 1.5625

and given d.f1 = 24 and d.f2 = 15

We can find F(α/2) using excel function =FINV( α/2, d.f1 , d.f2 )

and  F1-(α/2) using excel function =FINV( 1- α/2, d.f1 , d.f2 )

We are given confidence level = 0.90 ; therefore α = 1 - confidence level = 1 - 0.90 =0.10

So α = 0.10 , α/2 = 0.05 and 1 - (α/2) = 0.95

For F(α/2)     =FINV( 0.05,24,15 ) = 2.2878

For F1-(α/2)     =FINV( 0.95,24,15 ) = 0.4745

Lower bound =    Fo/2 = 1.5625 2,2878

Lower bound = 0.6830

Upper bound =  2 F1-(a/2) = 1.5625 0.4745

Upper bound = 3.2932

Therefore 90% confidence interval for the ratio of the two population variances associated with the two tests is

( 0.6830 , 3.2932 )